A cooler is 60 cm long, 24 cm wide, and 30 cm tall. What is the capacity in liters?
A. 4.32
B. 43.2
C. 432
D. 43,200


Thank you!

Answer by jim_thompson5910(35256)

Volume:

V = LWH

V = 60*24*30

V = 43200


So the volume is 43200 cubic centimeters.


Let's convert that to liters


1 L = 1000 cm^3

1/1000 L = 1 cm^3

1 cm^3 = 0.001 L

43200*1 cm^3 = 43200*0.001 L

43200 cm^3 = 43.2 L


So the volume of the cooler is 43.2 liters


So the answer is choice B